HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

DOUBLE RESONANCE EXCITATION OF OVERTONES AND COMBINATION BANDS IN JET-COOLED N_{2}O

\underline{J. BOSMANN,} \underline{F. PETERS} , \underline{K. VAVRA}, \underline{G. W. FUCHS} , \underline{T. F. GIESEN} , Laborastrophysik, Universität Kassel, Heinrich-Plett-Straße 40, 34132 Kassel,

Germany

Propylene oxide (PO) is one of the most discussed chiral molecules since its astronomical detection ^{1}. For future applications such as chiral detection of PO by three-wave mixing techniques we would like to aquire its ro-vibrationally resolved spectrum. In our investigations we found the ro-vibrationally resolved spectrum of PO to be quite complex and thus the assignment of ro-vibrational transitions relying solely on an IR cavity ring-down spectrometer not feasible.

In a first step we present the proof of principle concept of combining an IR cavity ring-down (2.7-4.0 µm) and chirped-pulse Fourier transform millimeter wave (100-112 GHz) spectrometer to excite double resonance processes in jet-cooled N_{2}O. In our setup the three components IR, MW and molecular jet are each aligned along the axis of a cartesian coordinate system and both excitation sources are linearly polarized, the polarization angle of the IR source is freely adjustable. We observed rotational transitions in the vibrational ground state (00^{0}0), the first overtone of the symmetric stretching mode (20^{0}0) around 3.9 µm and the combination band of symmetric and asymmetric stretching mode (10^{0}1) around 2.875 µm^{2}. Next a racemic mixture and pure enantiomers of PO are probed with the same setup. The double resonance procedure allows for precise assignment of the ro-vibrationally resolved spectrum. Results will be compared to theoretical predictions.

1doi:1606.07483, B. A. McGuire et al., ”Discovery of the Interstellar Chiral Molecule Propylene Oxide (CH3CHCH2O)”,Science 352 (2016)

2doi:10.1021/ed085p121, M. Bryant, S. Reeve, W. Burns, “Observation and Analysis of N2O Rotation Vibration Spectra”, J. Chem. Educ., 85 (2008)

p-number: p204  Submitted on Thu, 15 Jun 2023 22:01:33 +0200